# CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning

### v1

@@ 4.2.4 The Importance of Scalable PMP
A key benefit of using code data for PMP is the vast availability of publicly accessible, high-quality code-preference pairs, ensuring diversity. To validate scalability, we vary the number of training pairs for CodePMP and retrain models with different amounts of data. As shown in Figure 5, overall, increasing the number of code-preference pairs consistently improves BoN accuracy in both mathematical and logical reasoning tasks across model sizes, with no sign of diminishing returns. This indicates that further scaling the code-preference data would likely yield additional performance gains, underscoring the importance of building a scalable PMP pipeline.
(a) GSM8K / 1.5B
(b) MATH / 1.5B
(c) Reclor / 1.5B
(d) LogiQA2.0 / 1.5B
(e) GSM8K / 7B
(f) MATH / 7B
(g) Reclor / 7B
(h) LogiQA2.0 / 7B
@@
Figure 5: Increasing the number of code-preference pairs consistently improves Best-of-N accuracy in both mathematical and logical reasoning tasks across model sizes, with no evident signs of diminishing returns. Note that the horizontal axis is scaled by $\sqrt{2}$, and the gray dashed line represents the results without CodePMP.
@@ 5 Ablation Studies
In this section, we present a detailed analysis of CodePMP design. Unless otherwise stated, all experiments used the 1B model due to resource limitations and present the results of mathematical reasoning due to page limitation. Results of logical reasoning refers to Appendix E.2.
@@ 5.1 Impact of Pair Construction
Model-Generated Pairs Comparison
@@
Web-Crawled vs GitHub-Sourced Pairs
We also compare GitHub-sourced code with web-crawled data( Askell et al. (2021)) from platforms such as StackExchange and Reddit. As shown in Figure 7(a), GitHub-sourced pairs (“Source Code”) consistently outperform those from web platforms (“Webpage”), particularly as the number of solutions (N) increases. Moreover, the performance improvement of GitHub-sourced pairs shows no sign of plateauing, highlighting the importance of diverse, high-quality source code in building a scalable PMP pipeline.
@@
Figure 6: Comparisons of BoN accuracy across different construction methods and pair sources.
(a) With or without EOC token.
(b) Using WSD or WCD learning rate schedulers.
Figure 7: Comparisons of BoN accuracy for different settings: EOC token and lr schedulers.
5.2 Impact of EOC Token
Experiments by( Askell et al. (2021)) show that adding an end-of-context (EOC) token to each sequence significantly improves overall performance. To explore its impact in the context of CodePMP, we compared performance with and without the EOC token. As shown in Figure 7(a), the EOC setting (“w/ EOC”) consistently underperform the setting without EOC tokens (“w/o EOC”) across different test tasks, which is opposed to ( Askell et al. (2021)) We attribute this discrepancy to the different model, data and evaluation settings.
5.3 Impact of Learning Rate Schedulers
In the CodePMP experiments, we use the warmup-stable-decay (WSD) learning rate scheduler( Hu et al. (2024)), which can effectively reduce the time required for scaling related experiments. Previous studies mainly employ a learning rate schedule with linear warmup followed by cosine decay, known as warmup-cosine decay (WCD). We compare the performance of WSD and WCD, as shown in Table 7(b), both schedulers yield similar results. Thus, to improve computational efficiency, we adopt the WSD scheduler for all experiments.
@@ 5.3 Impact of Learning Rate Schedulers
Figure 8: CodePMP improves sample efficiency and Best-of-N performance of the Gemma-2B reward models on reasoning tasks, highlighting its broad applicability across diverse LLM architectures.
Table 2: Performance gains on coding RM and general RM (RMBench) evaluations show that CodePMP not only improves reasoning tasks but also generalizes well across various tasks.
RMBench
Model PMP Coding Summary Chat Chat Hard Safety Reasoning
1.5B 0.6841 0.4154 0.4804 0.5351 0.3665 0.2751
0.758 0.6126 0.9050 0.4364 0.3698 0.6041
7B 0.6912 0.5839 0.4972 0.5022 0.5240 0.6804
0.7619 0.7668 0.9413 0.5373 0.4906 0.9116
5.4 Validating CodePMP on Other LLMs
To further evaluate the generalizability of CodePMP, we validate its performance on the widely adopted Gemma-2B model (Team et al., 2024). As illustrated in Figure 8, the application of CodePMP results in significant performance gains in both mathemantical reasoning and logical reasoning evaluations. This not only underscores the robustness of CodePMP but also demonstrates its broad applicability in improving sample efficiency and overall performance across diverse LLM architectures.
5.5 Performance on Coding and General RM Benchmarks
@@
We evaluate CodePMP on both code-specific and general reward modeling benchmarks. The CodeUltraFeedback_binarized test set serves as an in-domain evaluation, while RMBench provides an out-of-domain assessment. As shown in Table 2, models finetuned with CodePMP initialization consistently outperform those without CodePMP across various model sizes. These results demonstrate that CodePMP not only enhances performance in reasoning tasks but also generalizes well across a range of RM benchmarks.
@@ 6 Related Works
In the context of RLHF, reward models (RMs) have traditionally employed ranking models like Bradley-Terry and Plackett-Luce to represent human preferences (Bradley & Terry, 1952; Plackett, 1975; Cobbe et al., 2021b; Saunders et al., 2022; Lightman et al., 2023b; Wang et al., 2023b; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024; Stiennon et al., 2020; Nakano et al., 2021). More recently, probability-based approaches (Zhao et al., 2023; Jiang et al., 2023) have emerged, offering more precise predictions. Additionally, models such as Critique-out-Loud (Ankner et al., 2024) enhance RMs by integrating natural language feedback. Generative reward models (GRMs) (Yang et al., 2024b) further boost sample efficiency. Preference Modeling Pretraining (PMP) (Askell et al., 2021) introduces a novel pretraining phase, utilizing large-scale pairwise ranking data to enhance RM performance. Despite these advancements, many methods are hindered by the reliance on expensive manual annotations or limited datasets, constraining scalability. CodePMP mitigates this by automating preference data generation from code, significantly improving RM sample efficiency and reducing dependency on manual data collection.
@@
The inclusion of code in LLM pretraining has led to marked improvements in tasks such as commonsense reasoning (Madaan et al., 2022) and mathematical problem-solving (Liang et al., 2022; Shao et al., 2024; Yang et al., 2024a). Furthermore, code enhances general reasoning capabilities (Muennighoff et al., 2023; Fu & Khot, 2022; Ma et al., 2023). Recent studies (Dong et al., 2023; Ma et al., 2023) indicate that incorporating code during supervised finetuning strengthens LLMs, particularly in complex decision-making tasks. CodePMP takes a pioneering approach by utilizing scalable, synthetically generated code preference pairs, reducing the dependence on manual annotation (Dubey et al., 2024; Gemini-Team et al., 2024; Groeneveld et al., 2024; Bi et al., 2024). This methodology enhances sample efficiency and scalability in reasoning-intensive tasks, presenting new opportunities for further improving LLM performance.
@@
Improving reasoning capabilities in LLMs remains a significant challenge, with various advanced methods being proposed. Chain of Thought (CoT) prompting (Wei et al., 2022; Fu et al., 2023) improves reasoning by generating intermediate steps, while CoT combined with supervised finetuning (SFT) further enhances performance (Cobbe et al., 2021a; Liu et al., 2024; Yu et al., 2023). Other approaches focus on expanding inference time computation, such as problem decomposition (Zhou et al., 2022), search-based methods like MCTS (Xu, 2023), and using LLMs as verifiers (Huang et al., 2022; Luo et al., 2023). Reward models, including outcome-based (ORM) and process-based (PRM), have also shown success, with PRM delivering superior results (Lightman et al., 2023a; Wang et al., 2023b). Distinct from these approaches, CodePMP introduces a scalable preference model pretraining phase that can integrate seamlessly with all the aforementioned techniques.
7 Conclusion & Future Works
This paper introduces CodePMP, a scalable pretraining approach that leverages code-preference pairs to improve reasoning capabilities in large language models. Experimental results validate that CodePMP significantly improves sample efficiency and boosts performance on reasoning tasks.
For future work, we aim to extend CodePMP in two key directions. CodePrMP will focus on utilizing compiler and interpreter verifiability to provide low-cost process supervision signals. GenPMP will explore how to improve sample efficiency and the performance of generative reward models by integrating code data.

### v2

@@ 4.2.4 Scalability Analysis
A key benefit of using code data for PMP is the vast availability of publicly accessible, high-quality code-preference pairs, ensuring diversity. To validate scalability, we vary the number of training pairs for CodePMP and retrain models with different amounts of data. As shown in Figure 5, increasing the number of code-preference pairs consistently improves BoN accuracy in both mathematical and logical reasoning tasks across model sizes, with no sign of diminishing returns. This indicates that further scaling the code-preference data would likely yield additional performance gains, underscoring the importance of building a scalable PMP pipeline.
(a) GSM8K / 7B
(b) MATH / 7B
(c) Reclor / 7B
(d) LogiQA2.0 / 7B
Figure 4: Sample efficiency comparison for 7B models: CodePMP-initialized reward models achieve higher Best-of-N accuracy with the equivalent sample sizes, showing better data efficiency. Horizontal axis scales by $\sqrt{2}$. Green: with CodePMP; Blue: without CodePMP.
(a) GSM8K / 7B
(b) MATH / 7B
(c) Reclor / 7B
(d) LogiQA2.0 / 7B
@@
Figure 5: Scaling analysis of CodePMP for 7B models: more code-preference pairs consistently improve Best-of-N accuracy across reasoning tasks without diminishing returns. Horizontal axis scales by $\sqrt{2}$; gray dashed lines show baseline performance without CodePMP.
@@ 5 Ablation Studies
This section presents a detailed analysis of CodePMP design. Unless otherwise stated, all experiments use the 1B model due to resource limitations and present the results of mathematical reasoning due to page limitation. More ablation studies refer to Appendix.
@@ 5.1 Impact of Pair Construction
GitHub-Sourced Pairs vs Web-Crawled
We compare GitHub-sourced code with web-crawled data (Askell et al., 2021) from platforms such as StackExchange and Reddit. As shown in Figure 6(b), GitHub-sourced pairs (“Source Code”) consistently outperform those from web platforms (“Webpage”), particularly as the number of solutions (N) increases. Moreover, the performance improvement of GitHub-sourced pairs shows no sign of plateauing, highlighting the importance of diverse, high-quality source code in building a scalable PMP pipeline.
Model Generated Data vs Human Data
@@
Figure 6: Comparison of BoN accuracy across construction methods and data sources, demonstrating benefits of model-based construction and GitHub code.
@@
Figure 7: Cross-architecture performance comparison: CodePMP enhances reasoning performance across different model families (Gemma2 and Llama3.2), showing broad applicability.
5.2 Impact of Loss Function
CodePMP integrates both Reward Modeling (RM) and Language Modeling (LM) loss components. To evaluate their contributions, we conducted experiments comparing three configurations: RM loss only, LM loss only, and the combined approach. As shown in Table 2, the combined loss function consistently outperforms single-loss variants across all Best-of-N evaluation settings, with particularly notable improvements on the challenging MATH dataset. This empirical evidence indicates a complementary relationship where RM loss enhances preference ranking while LM loss preserves general language capabilities, collectively yielding more robust reward model performance.
BoN RM Loss LM Loss RM + LM Loss
GSM8K
N=32 0.834 0.8317 0.8393
N=64 0.8362 0.8271 0.8453
N=128 0.8332 0.8309 0.8362
N=256 0.8271 0.8226 0.8484
MATH
N=32 0.344 0.376 0.418
N=64 0.358 0.376 0.424
N=128 0.366 0.354 0.434
N=256 0.362 0.372 0.41
Table 2: Loss function comparison.
5.3 Cross-Architecture Generalization
To assess CodePMP’s generalization capabilities beyond the Qwen architecture family, we evaluated its effectiveness with Gemma2 and Llama3.2 as PMP/RM backbones on GSM8K, MATH, Reclor, and LogiQA-v2 benchmarks. As shown in Figure 7, CodePMP: (1) Consistently enhances reasoning performance across all model families, and (2) Improves robustness at larger $N$ values, mitigating performance degradation observed in non-initialized models.These results demonstrate that CodePMP generalizes effectively across diverse model architectures, suggesting broad applicability of the approach.
5.4 Performance on Larger Backbone Model
To investigate CodePMP’s performance on larger model scales, we applied the technique to Qwen2-72B. Table 3 presents results across mathematical and logical reasoning tasks.
BoN w/o PMP w/ PMP w/o PMP w/ PMP
GSM8K MATH
N=1 0.7718 0.7718 0.298 0.298
N=4 0.8453 0.8453 0.424 0.424
N=32 0.8529 0.8628 0.488 0.500
N=256 0.8249 0.8400 0.506 0.514
BoN Reclor LogiQA-v2
N=4 0.894 0.918 0.7117 0.7927
Table 3: Performance comparison on reasoning tasks for Qwen2-72B with and without CodePMP initialization. Note that only $N=4$ was tested for Reclor and LogiQA-v2.
Results show consistent improvements with CodePMP initialization across all benchmarks. Notably, performance gains increase with larger $N$ values on challenging tasks like MATH, indicating that CodePMP’s benefits scale effectively to larger model architectures. The significant improvement on logical reasoning tasks further demonstrates CodePMP’s scalability and broad applicability.
5.5 Performance on More Powerful Generator
To determine whether CodePMP maintains its effectiveness with more sophisticated generators, we conducted experiments with two advanced models: Qwen2-Math-7B-Instruct (specialized for mathematical reasoning) and Qwen2.5-32B-Instruct (a substantially larger general-purpose model).
BoN GSM8K MATH
w/o PMP w/ PMP w/o PMP w/ PMP
N=4 0.8544 0.8931 0.690 0.724
N=32 0.8446 0.8795 0.643 0.698
N=256 0.8256 0.8590 0.614 0.690
Table 4: BoN accuracy with specialized mathematical generator (Qwen2-Math-7B-Instruct).
BoN GSM8K MATH
w/o PMP w/ PMP w/o PMP w/ PMP
N=4 0.9604 0.9688 0.798 0.820
N=32 0.9573 0.9581 0.768 0.792
N=256 0.9566 0.9634 0.752 0.798
Table 5: BoN accuracy with large-scale generator (Qwen2.5-32B-Instruct).
Tables 4 and 5 demonstrate that CodePMP’s benefits persist across different generator architectures. With the specialized Qwen2-Math-7B-Instruct (Table 4), we observe substantial improvements on both GSM8K and MATH. These gains remain consistent with the much larger Qwen2.5-32B-Instruct model (Table 5), despite it being significantly larger than both the preference pair generation models (less than 7B parameters) and the reward model itself (Qwen2-7B).
These findings demonstrate that reward models trained on synthetic preference data from smaller models can effectively guide more powerful and specialized generators, confirming CodePMP’s robustness and cross-scale applicability. This is particularly significant as it suggests that relatively modest investments in reward model training can yield benefits even when deployed with state-of-the-art generation systems.
5.6 Performance on General RM Benchmarks
@@
We further evaluate CodePMP on general reward modeling benchmarks (RMBench) to assess its applicability beyond reasoning tasks. RMBench provides an out-of-domain assessment covering various tasks including summarization, chat quality, and safety. As shown in Table 6, models fine-tuned with PMP consistently outperform those without PMP across various model sizes and tasks.
Model PMP RMBench
Summary Chat Chat Hard Safety Reasoning
1.5B 0.4154 0.4804 0.5351 0.3665 0.2751
0.6126 0.9050 0.4364 0.3698 0.6041
7B 0.5839 0.4972 0.5022 0.5240 0.6804
0.7668 0.9413 0.5373 0.4906 0.9116
Table 6: Performance on RMBench shows that CodePMP generalizes well across various general LLM tasks.
These results demonstrate that CodePMP enhances performance not only in reasoning and coding tasks but also improves generalization across a broad range of RM benchmarks. These findings provide compelling evidence for CodePMP’s broad applicability across multiple domains beyond the reasoning tasks that were our primary focus.
@@ 6 Related Works
In the context of RLHF, reward models (RMs) have traditionally employed ranking models like Bradley-Terry and Plackett-Luce to represent human preferences (Bradley and Terry, 1952; Plackett, 1975; Cobbe et al., 2021; Saunders et al., 2022; Lightman et al., 2023; Wang et al., 2023b; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024; Stiennon et al., 2020; Nakano et al., 2021). More recently, probability-based approaches have emerged, offering more precise predictions. Additionally, models such as Critique-out-Loud (Ankner et al., 2024) enhance RMs by integrating natural language feedback. Generative reward models (GRMs) further boost sample efficiency. Preference Modeling Pretraining (PMP) (Askell et al., 2021) introduces a novel pretraining phase, utilizing large-scale pairwise ranking data to enhance RM performance. Despite these advancements, many methods are hindered by the reliance on expensive manual annotations or limited datasets, constraining scalability. CodePMP mitigates this by automating preference data generation from code, significantly improving RM sample efficiency and reducing dependency on manual data collection.
@@
The inclusion of code in LLM pretraining has led to marked improvements in tasks such as commonsense reasoning (Madaan et al., 2022) and mathematical problem-solving (Liang et al., 2022; Shao et al., 2024; Yang et al., 2024). Furthermore, code enhances general reasoning capabilities (Muennighoff et al., 2023; Fu et al., 2022; Ma et al., 2023). Recent studies (Dong et al., 2023; Ma et al., 2023) indicate that incorporating code during supervised finetuning strengthens LLMs, particularly in complex decision-making tasks. CodePMP takes a pioneering approach by utilizing scalable, synthetically generated code preference pairs, reducing the dependence on manual annotation (Dubey et al., 2024; Gemini-Team et al., 2024; Groeneveld et al., 2024; Bi et al., 2024). This methodology enhances sample efficiency and scalability in reasoning-intensive tasks, presenting new opportunities for further improving LLM performance.
@@
Improving reasoning capabilities in LLMs remains a significant challenge, with various advanced methods being proposed. Chain of Thought (CoT) prompting (Wei et al., 2022; Fu et al., 2023) improves reasoning by generating intermediate steps, while CoT combined with supervised finetuning (SFT) further enhances performance (Cobbe et al., 2021; Liu et al., 2024; Yu et al., 2023). Other approaches focus on expanding inference time computation, such as problem decomposition (Zhou et al., 2022), search-based methods like MCTS (Xu, 2023), and using LLMs as verifiers (Huang et al., 2022; Luo et al., 2023). Reward models, including outcome-based (ORM) and process-based (PRM), have also shown success, with PRM delivering superior results (Lightman et al., 2023; Wang et al., 2023b). Encouragingly, CodePMP introduces a scalable preference model pretraining phase that can integrate seamlessly with all the aforementioned techniques.
7 Conclusion and Future Works
We propose CodePMP, a scalable preference model pretraining method that leverages synthetic code-preference pairs to boost reasoning in large language models. Experiments demonstrate that CodePMP markedly enhances both sample efficiency and performance across diverse reasoning tasks, validating the effectiveness of code-based preference pretraining. Future directions include CodePrMP, which will utilize compiler/interpreter verifiability for low-cost process supervision, and GenPMP, aimed at improving generative reward models through code-based pretraining.
Limitations
Our current implementation has several limitations. First, the synthetic preference pairs rely on models with predetermined parameter sizes, potentially missing nuanced preference signals that more sophisticated approaches might capture. While we demonstrate broad applicability across model families, architectural differences may affect performance in ways not fully explored in this work. Our reliance on GitHub data introduces potential biases stemming from the composition of public repositories. Additionally, our evaluation focuses primarily on mathematical and logical reasoning, leaving the method’s effectiveness for other reasoning modalities (e.g., commonsense or causal reasoning) less thoroughly examined. Future work should address these limitations to further enhance the generalizability and robustness of the approach.
Ethics Statement
CodePMP introduces several important ethical considerations. By enhancing LLMs’ reasoning capabilities, it could significantly impact decision-making systems that affect human lives, necessitating careful deployment and monitoring. While we utilize publicly available code, we recognize the importance of intellectual property rights and have focused on data with permissive licenses. Our approach reduces reliance on human annotation, potentially mitigating certain biases while possibly introducing others derived from the training data or model preferences. These trade-offs require ongoing evaluation and refinement to ensure fair and beneficial applications. As with any technology that enhances AI capabilities, responsible deployment with appropriate safeguards is essential.